40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Feb 13 15:02:23 2011, kiwamu, Update, Green Locking, X arm beam offcentering has been measured misposition.png
    Reply  Tue Feb 15 03:10:37 2011, kiwamu, Update, ASC, IR beam alignment for Xarm : TRX reduction 
       Reply  Wed Feb 16 02:04:11 2011, kiwamu, Update, ASC, IR beam alignment 
Message ID: 4278     Entry time: Sun Feb 13 15:02:23 2011     Reply to this: 4295
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: X arm beam offcentering has been measured 

The amounts of the X arm's beam off-centering have been measured by the A2L technique.

So now we are able to start aligning the IR beam axis in a quantitative way.

 

(motivation)

 Since we saw big residual motions at 1 Hz, 16 Hz on both the green beat note signal and the IR PDH signal (see #4268 and #4211),

we are suspecting that these noise come from an angle to length coupling.

In order to minimize the angle to length coupling, one thing we can do is to bring the beam spots to the center of ITMX and ETMX more precisely.

To do it, we have to quantitatively know how well the beam spots are on the center of the optics. Therefore I started measuring the amount of the beam off-centering.

 

(method)

 The A2L technique was used to measure the off-centering with the real-time lockin system, which has been recently embedded in the real-time code by Joe (see #4265).

The idea is the same as Yuta did before (see #3863).

But this time the excitation signal from the real-time oscillator was injected directly to the coil matrix on either ITMX or ETMX, at 18.13 Hz with the amplitude of about 400 cnt.

When the IR laser stays locked to the X arm, the LSC feedback signal is demodulated with the oscillator signal.

This demodulated signal gives the amount of the off-centering.

For this purpose I modified Yuta's A2L script such that we can use it also for the X arm.

 

(results)

 I obtained the following values:

     - ETMX

         PIT  = -1.61 mm

         YAW =  -0.918 mm

    - ITMX

         PIT = -3.76 mm

        YAW = -2.24 mm

I used the same calibration factor as that of Koji calculated (see #3020) for MC, in order to convert the results from the coil gain to the off-centering.

These values are consistent with the spots appearing on the CCD monitors.

 misposition.png

ELOG V3.1.3-